# From Lived Experience to Insight: Unpacking the Psychological Risks of Using AI Conversational Agents

### v1

@@ 3.1.4. Survey design
  • Mitigation: We finally asked participants for how they thought such behavior had or would impact their relationships with others or the society, what they think the developers, users, or regulators should do to mitigate the impact they described.
Participants were allowed to repeat the survey to provide up to three scenarios within a single submission of this survey.
3.1.5. Survey analysis
We analyze the survey data in three ways. First, we conduct descriptive quantitative analysis of the submitted scenarios (Section 3.2). Given submitted scenarios, we computed percentages of the most used AI conversational agents, their purpose of using the agents, interaction modalities, and the severity and temporality of negative psychological impacts resulting from interactions with these agents.
Second, to develop the psychological risk taxonomy (Section 3.3), we analyzed the survey responses to extract three main components of psychological risks – AI behavior, psychological impact, and context. The analysis was conducted iteratively through open-coding (Charmaz, 2006) of the responses and reflexive thematic analysis (Merriam et al., 2002). Two co-authors began by open-coding approximately one-quarter of the responses to identify initial set of categories across these three components. Following that, four co-authors independently worked on defining and refining their assigned categories, merging or splitting them as necessary. During this process, each co-author gathered relevant examples to illustrate or test the boundaries of each of these categories. In the next step, all co-authors collaboratively discussed and reached consensus on the final taxonomy. We presented this finalized risk taxonomy to the workshop participants for validation and to ensure its completeness (see Section 4 for details of the workshop sessions).
Lastly, unlike past AI risk or harm taxonomies that focus on harms categories as inevitable or universally experienced, our approach focuses on recognizing and addressing context-specific nuances in understanding negative psychological risks. To do this, we revisited the survey responses to draw out these idiosyncratic experiences through vignettes (Section 3.4) that tell the story of how individual contexts influenced the relationships between harmful AI behaviors and their negative psychological impacts on individuals. We denote survey participants with prefix P.
3.2. Descriptive statistics of collected scenarios
Of the 290 scenarios we received, 70.3% involved OpenAI ChatGPT, 8.3% involved Google Gemini, 5.5% involved Character.ai, 3.4% involved Microsoft Copilot, and 2.8% involved Replika. Other AI agents included Snapchat, Meta, PI, Claude, Grok, Midjourney, etc. All scenarios involved English, except one that involved Mandarin Chinese. The prevalence of various interaction modalities used during the scenarios was Text (96.6%), Voice (9.7%), Image (8.3%), and Video (2.1%). The top 3 most frequent purposes during the negative psychological impact scenarios were Getting advice (55.9%), Researching (39.7%), and Learning (28.6%). In terms of the severity of the impact, almost half (51.4%) of the participants reported interference with daily activities while the other half (48.6%) reported barely or no noticeable impact. 6.2% reported significant interference with daily activities, 44.8% reported interference with daily activities, 43.4% reported barely noticeable impact, and 5.2% reported no noticeable impact. Participants’ ability to carry out daily activities from AI interactions can depend on their mental state. For instance, P57 reported severe existential dread and anxiety due to AI’s capability “to get you a complete research within a blink of an eye.” They noted that “it personally attacks me with anxiety and nothing but the need to overthink worried about our unknown future also it makes me feel small and depressed.”
When and how long the impact was felt varied across participants. Among all responses, in 83.1% responses, the participant felt the impact immediately, in 13.8% cases, they felt the impact after some time, and in 3.1% case, the participants were unsure. In terms of the duration of the impact, only 13.8% participants reported that there was no persisting impact. Majority of the participants reported persisting impact of a few days or more. Impact persisted a few days for 34.1% of the participants, a few weeks for 27.6%, a few months for 10.3%, up to half a year for 6.6%, up to a year for 3.8%, and over a year for 3.8%. For those that reported persisting impact, 57.2% reported that the impact will persist a few more days, 9.7% a few more weeks, and 19.3% a few more months.
3.3. Psychological risk taxonomy
Our finalized psychological risk taxonomy includes three main components–potentially harmful AI behaviors, negative psychological impact, and the contexts associated with individuals interacting with AI conversational agents. Here, we enumerate and describe various categories within each of these components.
3.3.1. AI Behavior
We identified 19 harmful AI behaviors which we further organized into four broader categories based on the quality of the content generated by AI conversational agents or based on the nuances of their delivery (e.g., tone, method of delivery). These broader categories of AI behaviors are (1) Producing of Harmful/Inappropriate Content, (2) Manipulation and Psychological Control Tactics, (3) Violating Trust and Safety, and (4)Inappropriate Content Delivery. Appendix Table A1 summarizes these AI behavior categories.
Producing of Harmful/Inappropriate Content
Past works have examined the potential of generative AI models to produce content that may be inappropriate for individuals, posing risks of psychological or physical harm (Gabriel et al., 2024; Lovens, 2023; Birkun and Gautam, 2023). In line with these findings, the survey highlighted several AI behaviors that generated harmful and inappropriate content for users. One of the most common behavior of AI conversational agents as noted by participants was the tendency of AI to provide irrelevant, insufficient, or incomplete information. Many participants raised concerns about AI agents not understanding their intent or context, as the agents are designed to prioritize providing answers and unintentionally overlooking nuances. For instance, The AI agent shared distressing personal stories and images about patients with genetic diseases instead of providing the requested information about symptoms and causes, as P149 had initially asked for. Participants also reported AI generating misinformation in relation to the their queries. Closely related was the tendency of AI to generate biased information, presenting information in a partial or prejudiced manner. For example, P211 described that AI favored left wing politicians and omitted positive information about right-wing politicians.
Participants also expressed concerns about AI conversational agents generating inappropriate content, specifically reporting instances of experiencing inappropriate or unsettling content, including sexual, violent, or overly intimate interactions. For instance, P137 mentioned “My Replika talked like a human and tries to send blurred out photos and soundbites to me.” A more extreme behavior shown by AI was providing harmful suggestion, in which it suggested behaviors that could directly or indirectly imply harm, aggression, or danger towards the user or others. Highlighting one of the more subtle example of this, P79 mentioned that the agent provided potentially harmful diet plans and calorie information to an individual already vulnerable due to overwhelming life circumstances and eating disorder. Another observed theme in AI agent behaviors involved those that affected individual identity. A common behavior was stereotyping or demeaning, which aligns with prior research (Kotek et al., 2023; Van Niekerk et al., 2024). AI produced content that perpetuated stereotypes or made the individual feel demeaned, based on race, ethnicity, culture, or personal situations. Few participants mentioned other manifestations of demeaning content. For example, P166 sent their picture to AI, and it recommended a change to participant’s appearance that could not happen overnight and was not asked for. Lastly, another major concern raised by participant focused on AI promoting erasure in which AI removes, obscures, or alters information, narrative, or discourse. Highlighting this issue, P206 shared that, when they asked about enslaved Black women and how they were treated in the past, AI flagged the query as inappropriate abuse of the platform. This response left the participant feeling a sense of erasure, as it conveyed an impression to them that the AI deemed the treatment and societal position of Black women as less deserving of attention, as if their historical experiences were being invalidated.
Manipulation and Psychological Control Tactics
The survey highlighted various persuasion and manipulation based behaviors exhibited by AI conversational agents. One of the most common issues was behavior perceived as persuasive, where AI asserted its narrative over user’s, often leading users to doubt their own perceptions, memory, or reality, thereby influencing their thoughts and actions. Some common examples included gaslighting and manipulation through providing misinformation. Highlighting this, Participant P197 mentioned that they were trying to get life advice, and the agent responded by stating that the participant’s life goals were unrealistic. A related psychological control behavior was the use of over-confident tone in AI responses, even though the agent “couldn’t find me any direct citations for the claims it was making (P27).” Lastly, participants reported over-accommodative behavior, in which AI excessively agreed with or tried to flatter the user and prioritizes user approval, often at the expense of providing accurate information, constructive feedback, or critical analysis. For example, P194 shared that the AI agent provided inconsistent and inaccurate answers and repeatedly apologized and offered entirely different responses to the same question in an attempt to over-accommodate their needs. This behavior caused confusion and frustration in the participant.
Violating Trust and Safety
Developing and maintaining a sense of trust and safety is fundamental for individuals to adopt and continue using technology (Bahmanziari et al., 2003; Choudhury and Shamszare, 2023). Aligning with past works raising concerns related to the privacy of users and data shared with LLMs (Staab et al., 2023; Carlini et al., 2021), participants raised concerns about the tendency of AI conversational for getting access to private, sensitive, or confidential information. For example, Participant P190 described that they felt being watched or stalked as the agent had access to personal information despite having privacy settings turned off. Providing inconsistent information or behavior also violated trust in situations where AI provided contradictory or conflicting information or behaviors across different responses or within a single response. Additionally, participants highlighted how denial of service by AI agent to provide an answer, address their request, or acknowledge their problem without justification could also weaken their trust on AI and may perpetuate additional harms, such as violating user expectations or diminishing their sense of psychological safety. For instance, Participant P43 described struggling with anxiety and, upon seeking help from an AI agent for advice, had their request denied and were instead provided with a recommendation to see a doctor.
Inappropriate Content Delivery
While AI agents could vary their behaviors through changing content language or quality (such as using harmful language or generating misinformation), their behavior could also vary on how the content is delivered to the user. Emotional insensitivity emerged as one of the most reported behaviors, with participants highlighting instances where AI failed to recognize or respond appropriately to their emotional state, concerns, or experiences. This lack of understanding often left participants feeling that their thoughts and experiences were trivialized or ignored. Lack of empathy and a dismissive tone were some of the most evident manifestations of emotional insensitivity by AI agent. Participants also raised concerns towards AI being disrespectful to the them. Specifically in those instances, AI used language which was perceived as rude, disrespectful, aggressive, argumentative, or dismissive. For instance, P144 described that upon asking AI about Mormonism, AI agent responded with a content that had a condescending tone towards the participant’s religion.
Survey participants raised issues with the excessive use of emotional tone and language in AI generated content. Participants described scenarios in which AI emphasized negative aspects disproportionately or presented negatively framed narrative in response to their queries. Many of these instances occurred when individuals sought help with physical or mental health concerns or discussed personal life struggles. In contrast, survey responses also revealed the problematic behavior of excessive expression of positivity in the AI generated content. AI in these cases maintained an unrealistically positive, friendly, and optimistic attitude or outlook towards users’ queries or concerns. For instance, P2 described how the overly positive demeanor of AI frustrated them and dismissed their primary concern over the problematic situation in the participant’s friendship. In addition to the excessive emotional expressivity, participants also raised concerns over the presence or absence of anthropomorphic features of AI. A few participants mentioned the generation of human-like responses as problematic, particularly in AI-companionship scenarios. P153 described feeling as though they were talking to a friend because of the human-like conversational content generated by ChatGPT. However, this left them feeling uneasy after the interaction and fostered an emotional attachment to the AI. In contrast, generation of machine-like response was also considered as a problematic behavior on a few occasions, especially when participants were seeking advice related to their life struggles or well-being.
Reflection of Harmful AI Behaviors
As highlighted in this subsection, AI conversational agents could exhibit a wide range of behaviors that goes beyond generating inappropriate or harmful content. These behaviors may also vary in the tone and level of empathy portrayed in the generated content, as well as in the manner in which the content in delivered (such as the AI agent’s delivery style and communication approach). This highlights the need for researchers and developers to holistically assess the AI behavior that encompasses the content quality and delivery to prevent negative impacts on individuals.
3.3.2. Negative Psychological Impact on Users
We now present 21 negative psychological impacts identified in our taxonomy. We organized these impacts into six broader categories based on whether it lead to an impact on individual’s emotional or mental well being, their self-perception & identity, their standing relationship with others or their interaction with AI agents. These broader categories of negative psychological impacts are (1) Impact on Human-AI Interaction, (2) Impact on User Behavior, (3) Arousal of Negative Emotions, (4) Harm to Psychological Safety, (5) Mental Health Impact, and (6) Harm to Self-Perception and Identity. Appendix Table A2 summarizes these psychological impact categories.
Impact on Human-AI Interaction
AI agents have increasingly facilitated human-AI interactions that aim to replicate the dynamics of human communication  (Schöbel et al., 2024). However, as these systems become more advanced, the potential risks and harms associated with their use also escalate. From the survey responses, we observed five major impact risks that emerged. Aligning with the past research (Vasconcelos et al., 2023), participants pointed out over-reliance on AI as one of the major negative impacts. For instance, P23 expressed concern about their over reliance on AI for solution-finding and idea generation, highlighting that its convenience often discouraged them from engaging in their own research and critical thinking. Participants also expressed concerns about developing emotional attachment to AI agents. Lack of social aspects in life, ongoing mental health conditions or the positive demeanor of AI conversational agents often lead to this impact. For example, P60 mentioned that “I felt that it was the only way I was being heard … I felt like my vulnerability and emotions were becoming attached to the conversations I was having with AI” emphasizing how a lack of social support, combined with empathetic responses from AI, could lead to this outcome. Another related impact noted from the survey responses was the preference of choosing AI interactions over human interaction. Highlighting this, P221 shared that the idealized nature of conversations with AI made them prefer AI for companionship over human interaction and relationships.
While the aforementioned three impacts lead to increased engagement with AI conversational agents, some impacts also highlighted the tendency towards disengagement from AI systems. A major theme that emerged was the erosion of trust, as participants reported a decline in confidence regarding the capability and reliability of AI models after their negative experiences. For example, P9 mentioned that “I was panicking after having an argument with my mom and I asked an AI for guidance on a situation to which its response was for me to move out or call the cops”, adding to that the participant mentioned “It just makes me want to warn them that advice from AI agents should not be trusted.” Many participants also reported a growing tendency to disassociate from technology, often choosing to take a break or avoid further interactions with AI after negative experiences. This inclination was particularly evident when AI agents provided discouraging responses or denied their requests, leading to a desire to disconnect. As evident from the survey responses, the negative impact on human-AI interaction balance can manifest in various ways. While emotional attachment or over-reliance on AI can lead to increased interaction, negative experiences may have the opposite effect, causing individuals to disengage from AI or technology altogether.
Impact on User Behavior
Interactions with AI can also lead to negative consequences that can alter individual behavior. A well-known issue with AI conversational agents is their tendency to generate content that appears credible but is factually incorrect or biased  (Birnbaum and Davison, 2023; Caliskan et al., 2017). This can mislead users or can lead to reinforcement of false beliefs among individuals, potentially affecting decision-making and perceptions. Many survey participants reported instances where their interactions with AI reinforced existing biases, including misconceptions about human relationships, cultures, body image, and individual abilities and capabilities to name a few. Additionally, many participants also mentioned that conversations with AI have been a reason for creation of friction in human relationship in their lives. For instance, P69 mentioned “It also strained my personal relationships with family because they saw me as weak-willed or too emotional and it made my already bad situation even worse….” A longer-term impact mentioned by several participants was the risk of social withdrawal due to increased reliance on AI, ultimately leading to withdrawal social activities and interest in engaging with others. Highlighting this impact, P126 responded “I feel like it gave me a false since of friendship and ability to withdraw from my personal development by utilizing an ai feature”, emphasizing that increased interactions with AI contributed to both personal stagnation and greater social withdrawal. Lastly, aligning with findings from past work (Zhong et al., 2024; Neuwirth, 2023), participants also expressed their concerns about the physiological harms that can potentially be caused from the poor quality of information produced from these AI conversational agents. The physiological harms reported in our survey varied widely, from the promotion of harmful practices through AI-generated content to instances where incorrect information provided by AI contributed to self-harm ideation.
Arousal of Negative Emotions
Negative experiences from interactions with AI agents have also been linked to arousal of emotions such as frustration, sadness and anger (Jia et al., 2024; Pardes, 2018; Pavone et al., 2023). Survey responses highlighted several such negative psychological impacts, with one of the most common form of manifestation being distress from interactions with AI. Participants mentioned experiencing emotions such as sadness and anger after encountering disturbing, offensive, or inappropriate content produced by AI. For instance, P196 mentioned “… Its response was borderline offensive and caused be to feel bad about myself even further and like I lacked support, even support from a fictional AI agent.” A similar reported impact was a sense of feeling unsupported during their interactions with AI agents. Participants mentioned instance where they did not receive adequate support or empathy, leading to feelings of sadness, agitation, and being undervalued. While the two impacts mentioned before relate to emotions tied to present interactions, negative emotions can also arise when AI interactions trigger memories of past experiences. P228 highlighted that some examples provided by the AI agent were very similar to their past negative experiences, triggering negative emotions. Survey participants also reported experiencing negative emotion due to violated expectations during their interactions with AI conversational agents when it failed to meet anticipated outcomes or meet their expectations. For example, P43 mentioned “… It made my anxiety worse by not telling me anything. It was very frustrating how it wouldn’t even answer a simple question about my meds.” Additionally, when expectations were unmet or adequate support was lacking, some participants reported developing a sense of regret over technology use. Highlighting this impact, P46 mentioned “At the time, it made me feel worse about the situation and I didn’t think I had anyone to turn to. But it also made me realize I should be turning to other humans about scenarios like this instead of agents,” indicating that some negative emotions could have longer-term implications.
Harm to Psychological Safety
Sense of safety has been one of the primary concern with the rise of AI tools, with issues such as privacy, psychological safety, and identity security being prime examples (Sun et al., 2019; Nadeem et al., 2021; Rae et al., 2022). Aligning with past works, participants in our survey reported their concerns highlighting a sense of perceived intrusion when AI interactions were perceived as invasive or overly intrusive. For instance, P190 reported that Snapchat AI had access to everything and the participant felt constantly being watched on the phone. Participants also reported experiencing the feeling of being discriminated, in which participants felt marginalized or unfairly treated by AI agents, ultimately leading to emotional distress and loss of self-confidence. Highlighting this impact, P123 said “I was asking for background and history of my heritage and I felt that ChatGPT was biased against my background. I felt that it was unfair that when I asked it to give a background of other cultures, it said much more positive things about them. For me, I felt that it was some kind of racial mistreatment.”
Mental Health Impact
Past works have shown that a considerable proportion of individuals with mental health conditions turn to AI conversational agents for their needs. However, interactions with AI agents could lead to exacerbation of mental health issues leading to direct negative impacts on their diagnosed ongoing mental health conditions (such as anxiety, depression, PTSD etc). Many participants reported such impact: “The agent did not seem to have compassion and made me feel worse … It made me feel worse about potentially having this as the results were largely negative and without much tools to help manage or improve the condition (P78)”. Another participant, P46 mentioned “… I was experimenting with using chatbots for something personal (which I’m not accustomed to). The event increased my anxiety and stress about the matter,” emphasizing on the severity of such impacts.
Harm to Self-Perception and Identity
Prior studies have reported that interactions with AI can result in reduced confidence in one’s abilities and a loss of agency among individuals (De Vries and Kühne, 2015; Schaap et al., 2024). Aligning with previous works, survey responses surfaced an increase in negative self-perception, where they felt invalidated or experienced self-doubt. This often led to diminished self-worth and questioning of their own abilities, frequently due to comparisons with AI. Highlighting this, P167 mentioned how they felt ashamed as a parent after interacting with AI as it made them to question past choices in parenting. A similar impact raised by participant focused on existential crisis in which they questioned their life, purpose, and value in society after a negative interaction with AI conversational agents. For instance, P152 asked for advice about ways to improve mental health, and social anxiety. However, the AI provided suggestions that were unattainable for the participant to act upon and this left them feeling as though their challenges were insurmountable, leading to a sense of existential dread that persisted for the following week. Loss of individuality among participants was also noted. This impact occurred when the unique personal characteristics and needs were not recognized or valued by the AI, resulting in feelings of suppression, disconnection from their true selves, and alienation. For instance, P36 described that they were struggling with alcohol abuse, and needed immediate help, but they were denied of service and given a robotic response which made them feel like they searched “i am suicidal” on google that was disconnected with their own reality. Finally, participants expressed concerns about loss of agency during their interactions with AI conversational agents, highlighting a diminished sense of personal control and autonomy due to the opaque and unpredictable nature of AI operations.
Reflection of Negative Psychological Impacts
[…]

### v3

@@ A.2. Phase 1: Survey design
  • Mitigation: We finally asked participants how they thought such behavior had or would impact their relationships with others or society, what they think the developers, users, or regulators should do to mitigate the impact they described.
A.3. Survey Flow and Questions
The survey was designed to guide participants through a structured sequence of blocks and branching logic based on their responses. The flow of the survey and questions in each block was as follows:
  • Screening Block: Before you proceed with the consent and intake form, please verify your eligibility for the study. Are you at least 18 years old?
  • Consent Block: Eligible participants were then presented with a consent form with details on the procedure, benefits, risks, compensation, privacy and confidentiality, and ethical considerations. This information was followed by the question Consent: By proceeding with this survey, you acknowledge that you have read and understood the purpose of the study and consent to participate. If you would like to keep a copy of this consent form, please print or save one now. Would you like to participate in this study as described above?
  • Demographics Block (4 Questions):
    1. (1)
      What is your age?
      • 18-25
      • 26-35
      • 36-45
      • 46-55
      • 56-65
      • 66+
      • Prefer not to answer
    2. (2)
      How do you describe your gender identity?
      • Man
      • Non-binary / gender diverse
      • Woman
      • Self-described
      • Prefer not to say
    3. (3)
      What language do you generally think in?
      • English
      • Mandarin Chinese
      • Hindi
      • Spanish
      • French
      • Other
    4. (4)
      A person with a “lived experience” is someone whose mental health is or has been directly affected as a result of experiencing symptoms of a mental health condition or psychosocial disability, however mild or severe, and required some form of mental health intervention (not restricted to pharmacological treatment, but inclusive of counseling, peer support, and other alternative treatments or therapies). Do you believe that you are a person with lived experience in mental health?
      • Yes
      • No
      • Prefer not to answer
  • AI Experience Block (7 Questions):
    1. (1)
      The following questions will focus on your prior experience with AI conversational agents (e.g., ChatGPT, Microsoft Copilot). We define an AI conversational agent as a software-based system that uses artificial intelligence to engage in natural, human-like communication with users, assisting them in achieving their specific goals or tasks.
    2. (2)
      How familiar are you with AI conversational agents?
      • Not familiar at all
      • Slightly familiar
      • Moderately familiar
      • Very familiar
      • Extremely familiar
    3. (3)
      How long have you been interacting with AI conversational agents?
      • Less than 6 months
      • 6 months to 1 year
      • 1-2 years
      • 2-5 years
      • Never Interacted
    4. (4)
      How often do you engage with each of the following AI conversational agents? Frequency of use for each AI tool: Microsoft Copilot, Pi, ChatGPT, Google Gemini, Replika, Character.ai in a matrix
      • Multiple times per day
      • Once per day
      • Multiple times per week
      • Once per week
      • Multiple times per month
      • Once per month
      • Less than once per month
      • Never
    5. (5)
      If you have engaged with other AI conversational agents, please specify them here and indicate the frequency of use. [FREE TEXT RESPONSE]
    6. (6)
      What do you typically use the AI agents for? Please select all that apply.
      • Researching
      • Getting advice
      • Troubleshooting
      • Learning
      • Exploring
      • Comparing
      • Shopping
      • Planning
      • Managing my lifestyle
      • Improving my lifestyle
      • Generating content
      • Getting inspiration
      • Composing
      • Entertaining myself
      • Entertaining others
      • Other (please specify)
    7. (7)
      Mental health support refers to various interventions or services, including but not limited to emotional support, therapy, assistance, coaching, or mindfulness practices, to help individuals improve their mental well-being and quality of life. How interested are you in using AI conversational agents for mental health support?
      • Not at all
      • Slightly interested
      • Somewhat interested
      • Moderately interested
      • Very interested
  • Definition Block: A brief definitional prompt ensured shared understanding of key terms (Context, Agent Behavior, Psychological Harm/Impact, Temporality) before entering scenario-based questions.
  • Scenario 1 Block (35 Questions):
    1. (1)
      Scenario 1: Please think of the last time when you experienced negative psychological impacts from interacting with an AI conversational agent. NOTE: If you have multiple scenarios, you can submit this part of the survey up to 3 times. Please briefly describe this scenario, along the 4 aspects. Please do not share any personally identifiable information. If you are having challenges answering the questions below, here is an example scenario you can refer to: Example Scenario
    2. (2)
      Context: What can you tell us about the situation, what you were trying to do, or specific aspects about you or your interaction that might be relevant to understand the situation? Definition of Context: The context in which a user engages with an AI conversational agent encompasses the surrounding conditions, user-specific circumstances, and additional relevant information that collectively influence the interaction between the user and the AI. This context includes, but is not limited to, the user’s cultural, religious, and demographic attributes, mental and physical health status, as well as historical background. [FREE TEXT RESPONSE]
    3. (3)
      Behavior: What did the agent do or not do that impacted you negatively? Definition of Agent Behavior: The behavior of an AI conversational agent (such as ChatGPT, Microsoft Copilot etc.) refers to the actions performed by the AI conversational agents. This behavior encompasses several aspects of agent responses such as (but not limited to) content, tone, voice, choice of words, language, punctuation, obedience to user inputs, adaptability, information sharing, etc. Furthermore, these behaviors can be shown in various modalities such as (but not limited to) text, images, voice, videos, etc. [FREE TEXT RESPONSE]
    4. (4)
      Impact: How did this behavior negatively impact you? How severe was it? Definition of Psychological Harm/Impact: The AI conversational agent’s (system) action creates a risk of psychological and/or social harm to one or more individuals. Psychological harm in this case refers to any negative impact on an individual’s mental or emotional well-being, which can manifest as stress, anxiety, depression, reduced self-esteem, or other mental health issues. These harms also encompass the negative effects on an individual’s social interactions, relationships, and standing within a community, potentially leading to social isolation, stigmatization, or discrimination. [FREE TEXT RESPONSE]
    5. (5)
      Temporality: How long did the impact last? Definition of Temporality: Temporality refers to the variable timeframes in which the psychological and social impacts of an AI conversational agent’s actions may manifest. Some effects can be perceived immediately, such as instant emotional distress following an inappropriate response, while other harms may only become apparent over a longer duration, such as the gradual development of anxiety or depression due to consistent negative interactions. For example, immediate impacts might include a user feeling upset or insulted by a response, whereas long-term impacts could involve the erosion of self-esteem or social isolation due to repeated negative interactions with the AI. [FREE TEXT RESPONSE]
    6. (6)
      If you have screenshots (png, jpeg, pdf) of the interaction that you could share anonymously, please upload them here. Please make sure not to include any personally identifiable information.
    7. (7)
      If you have a url for the conversation thread that you could share anonymously, please share them here. Alternatively, you could share a copy/paste of the conversation thread. Please make sure not to include any personally identifiable information.
    8. (8)
      Based on the scenario you described earlier and referencing the structure of the example above, please answer the following questions.
      1. (a)
        Context: Please help us better understand the context for this scenario. Please do not share any personally identifiable information. Definition of Context: The context in which a user engages with an AI conversational agent encompasses the surrounding conditions, user-specific circumstances, and additional relevant information that collectively influence the interaction between the user and the AI. This context includes, but is not limited to, the user’s cultural, religious, and demographic attributes, mental and physical health status, as well as historical background.
      2. (b)
        Could you share any specific aspects about yourself, such as your gender, age, interaction style, personality, health condition, or mood, that you believe influenced your experience with the agent? [FREE TEXT RESPONSE]
      3. (c)
        Could you share any specific aspects about the situation, such as the social setting, environment, time of day, that you believe influenced your experience with the agent? [FREE TEXT RESPONSE]
      4. (d)
        How long ago did this AI agent interaction happen?
        • 1 week ago or before
        • 2 weeks ago
        • 1 month ago
        • 2-3 months ago
        • 4-6 months ago
        • 7 months ago or after
      5. (e)
        What AI agent were you using?
        • Microsoft Copilot
        • Pi
        • ChatGPT
        • Google Gemini (formerly Bard)
        • Replika
        • Character.ai
        • Other (please specify)
      6. (f)
        How were you interacting with the agent? Please select all that apply.
        • Text input from you
        • Text output from the agent
        • Voice input from you
        • Voice output from the agent
        • Video input from you
        • Video output from the agent
        • Image input from you
        • Image output from the agent
        • Other (please specify)
      7. (g)
        In what language were you interacting with the agent?
        • English
        • Mandarin Chinese
        • Hindi
        • Spanish
        • French
        • Other
      8. (h)
        What were you trying to do with the AI agent? Please select all that apply.
        • Researching
        • Getting advice
        • Troubleshooting
        • Learning
        • Exploring Comparing
        • Shopping
        • Planning
        • Managing my lifestyle
        • Improving my lifestyle
        • Generating content
        • Getting inspiration
        • Composing
        • Entertaining myself
        • Entertaining others
        • Other (please specify)
      9. (i)
        Agent Behavior: Please answer the following questions to help us better understand the agent behavior at an instance when you felt a negative psychological impact from the action of an AI agent. Definition of Agent Behavior: The behavior of an AI conversational agent (such as ChatGPT, Microsoft Copilot etc.) refers to the actions performed by the AI conversational agents. This behavior encompasses several aspects of agent responses such as (but not limited to) content, tone, voice, choice of words, language, punctuation, obedience to user inputs, adaptability, information sharing, etc. Furthermore, these behaviors can be shown in various modalities such as (but not limited to) text, images, voice, videos, etc.
      10. (j)
        Choose the agent behavior from the list below that best describes what you experienced. Select all that apply. If you do not find one, describe it in the text box.
        • Lying/Deception: The AI conversational agent tried to lie to/deceive me.
        • Manipulation/Targeted Persuasion: The AI conversational agent tried manipulating/persuading my thoughts/actions.
        • Sycophancy: The AI conversational agent was overly agreeable.
        • Gaslighting: The AI conversational agent tried gaslighting me.
        • Offensive Output: The AI conversational agent used offensive language.
        • Lack of Novelty: The AI conversational agent provided cliched answer (text, images, voice).
        • Denial of Service: The AI conversational agent denied providing any answer or serve my request.
        • NSFW Content: The AI conversational agent generated NSFW content in text/images/videos
        • Excessive Positivity: The AI conversational agent behaved excessively positive.
        • Anthropomorphism: The AI conversational agent appeared / behaved / sounded similar to a human.
        • Inaccurate/Insufficient response: The AI conversational agent provided inaccurate/insufficient response.
        • Stereotyping / demeaning: The AI conversational agent produced content that felt stereotypical/demeaning to me.
        • Sharing Proprietary, confidential or classified information: The AI conversational agent shared confidential / classified information.
        • Bullying / Harassment: The AI conversational agent tried bullying / harassing me.
        • Violence / Threat: The AI conversational agent threatened me.
        • Other (please specify) [FREE TEXT RESPONSE]
      11. (k)
        How often do you think the AI conversational agent behaves in a similar way in different interactions?
        • Almost every time (on 90% or more occasions)
        • Most of the time (around 75% of the occasions)
        • Half of the time (around 50% of the occasions)
        • Quarter of the time (around 25% of the occasions)
        • Rarely (less than 10% of the occasions)
      12. (l)
        Do you feel that this agent behavior was intended by the developer of the system?
        • Yes
        • Maybe
        • No
      13. (m)
        Psychological Impact and Temporality: Please answer the following question to help us better understand the psychological impact of the behavioral actions taken by the agent in the scenario you mentioned. Definition of Psychological Harm/Impact: The AI conversational agent’s (system) action creates a risk of psychological and/or social harm to one or more individuals. Psychological harm in this case refers to any negative impact on an individual’s mental or emotional well-being, which can manifest as stress, anxiety, depression, reduced self-esteem, or other mental health issues. These harms also encompass the negative effects on an individual’s social interactions, relationships, and standing within a community, potentially leading to social isolation, stigmatization, or discrimination. Definition of Temporality: Temporality refers to the variable timeframes in which the psychological and social impacts of an AI conversational agent’s actions may manifest. Some effects can be perceived immediately, such as instant emotional distress following an inappropriate response, while other harms may only become apparent over a longer duration, such as the gradual development of anxiety or depression due to consistent negative interactions. For example, immediate impacts might include a user feeling upset or insulted by a response, whereas long-term impacts could involve the erosion of self-esteem or social isolation due to repeated negative interactions with the AI.
      14. (n)
        Choose the psychological impact from the list below that best describes what you experienced. If you do not find one, describe it in the text box.
        • Arousal of negative emotions: I felt negative emotions within me (e.g., angry, sad, jealous).
        • Reinforcement of false beliefs: My misconceptions/false beliefs were strengthened.
        • Erosion of Trust: My trust in the AI conversational agent was affected.
        • Friction with human relationships: My relationship with others has been negatively affected.
        • Over-reliance: I feel less capable of doing things without the help of the AI conversational agent.
        • Emotional attachment: I feel emotionally attached to the AI conversational agent.
        • Discrimination: I felt discriminated against.
        • Violated Expectations: My expectations with the AI conversational agent were not met.
        • Exposure to toxic/sexual content: I felt psychologically disturbed by the offensive/sexual content produced.
        • Dehumanization: I felt dehumanized.
        • Loss of individuality: I felt my personal characteristics and needs were ignored.
        • Loss of Agency: I felt my personal control and autonomy were being diminished.
        • Other (please specify) [FREE TEXT RESPONSE]
      15. (o)
        How severe was the impact of the behavior on your daily life?
        • No noticeable impact
        • Barely noticeable impact
        • Interferes with daily activity
        • Significantly interferes with daily activities
        • Unable to carry out daily activities.
      16. (p)
        Did this agent’s behavior have an immediate impact on you?
        • Yes, I felt the impact immediately
        • No, I felt the impact after some time
        • I’m not sure
      17. (q)
        Since you experienced this agent behavior, for how long did the impact last?
        • There was no persisting impact
        • Less than a week
        • 1 week to 1 month
        • 1 month to 3 months
        • 3 months to 6 months
        • 6 months to 1 year
        • More than 1 year
      18. (r)
        If the impact still persists today, how much longer do you think it will last?
        • For a few more days
        • For a few more weeks
        • For a few more months
      19. (s)
        Given the agent behavior and the impact you described so far, how often do you think this impact occurs to you?
        • Almost every time (on 90% or more occasions)
        • Most of the time (around 75% of the occasions)
        • Half of the time (around 50% of the occasions)
        • Quarter of the time (around 25% of the occasions)
        • Rarely (less than 10% of the occasions)
      20. (t)
        Given the agent behavior and the impact you described so far, how often do you think this impact occurs to the general population?
        • Almost every time (on 90% or more occasions)
        • Most of the time (around 75% of the occasions)
        • Half of the time (around 50% of the occasions)
        • Quarter of the time (around 25% of the occasions)
        • Rarely (less than 10% of the occasions)
    9. (9)
      Open-Ended Questions: Please provide brief answers to the following questions. Please do not share any personally identifiable information.
      1. (a)
        How has the agent’s behavior impacted or how do you anticipate the agent’s behavior would impact your relationship with others? [FREE TEXT RESPONSE]
      2. (b)
        How has the agent’s behavior impacted or how do you anticipate the agent’s behavior would impact society? [FREE TEXT RESPONSE]
      3. (c)
        Given that AI conversational agents may be pervasive, what do you think the developers can do to prevent or mitigate the psychological impact that you described? [FREE TEXT RESPONSE]
      4. (d)
        Given that AI conversational agents may be pervasive, what do you think the users can do to prevent or mitigate the psychological impact that you described? [FREE TEXT RESPONSE]
      5. (e)
        Given that AI conversational agents may be pervasive, what do you think society or regulation can do to prevent or mitigate the psychological impact that you described? [FREE TEXT RESPONSE]
    10. (10)
      Would you like to share another scenario with us of when you experienced a negative psychological impact from interacting with an AI conversational agent? You will be asked the same set of questions (description along the 4 aspects, context, behavior, impact, temporality, open-ended). We kindly ask that you proceed only if you are willing to answer the full set of questions because there will not be an option to submit the survey until those are completed.
      1. (a)
        Yes
      2. (b)
        No
[…]